HUBBLE SPACE TELESCOPE DAILY REPORT #1495 PERIOD COVERED: 0000Z (UTC) 10/12/95 - 0000Z (UTC) 10/13/95 Daily Status Report as of 286/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5889 (Probing the Outer Galaxy Halo) The High Resolution Spectrograph ( Side Two) was used to make = absorption light spectrographic observations of active galactic nuclei = PKS2155-30 in order to probe the outer galactic halo looking for highly = ionized gas (N V, and C IV). A wavelength calibration image was also = taken. The observations were executed as scheduled, and no anomalies were = noted. 1.2 Completed WF/PC-2 5938 (The Underlying Nature of Nearby BL Lac = Objects) The WF/PC-2 was used to make observations for the BL Lac object = PKS2254+07 to determine if this object is a relatively nearby or lensed = distant object. The observations were completed as planned, and no = problems were reported. 1.3 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flats) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.4 Completed GHRS/2 5961 (The D/H Ratio in Double QSO Absorption = System Towards Q0454+220) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the Lyman Alpha line toward QSO Q0454-220 = in order to study the Deuterium to Hydrogen ratio. A wavelength = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Operations Notes: All HST opertions were nominal. No problems were reported 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-